OpenStudy (anonymous):

'Write an equation for the line that passes through (-5, -7) and is parallel to the line -5x+2y=27" ?How do I do it? Please someone walk me through this!

OpenStudy (amistre64):

parallel lines have the exact same equation except for the last part; so chop it off and reslove

OpenStudy (amistre64):

-5x + 2y = 27 -5x + 2y = N ; fill in the point and solve for N,

OpenStudy (amistre64):

-5(-5) + 2(-7) = N 25 - 14 = N 11 = N .................. parallel equation is: -5x + 2y = 11

OpenStudy (amistre64):



OpenStudy (anonymous):

Oh my goodnees! Thank you soo much!

OpenStudy (amistre64):

these parallel lines have the exact same basic equation; but differ only by a constant